What force must be exerted on the pedal cylinder of a hydraulic lift to support the weight of a 2411-kg car (a large car) resting on the wheel cylinder? The master cylinder has a 2.1-cm diameter and the second cylinder has a 29.2-cm diameter.


Round your answer to 2 decimal places


1
Expert's answer
2021-04-16T07:26:36-0400

The main equation is the following (see https://en.wikipedia.org/wiki/Hydraulic_press):


"\dfrac{F_m}{A_m} = \dfrac{F_2}{A_2}"

where "F_m" is the unknown force applied to the master cylinder, "A_m" is the area of the master cylinder, "F_2 = mg" is the force applied to the second cylinder ("m = 2411kg" is the mass of the car), and "A_2" is the area of the second cylinder.

The areas are:


"A_m = \pi d_m^2/4\\
A_2 = \pi d_2^2/4"

where "d_m = 2.1cm = 0.021m,\space d_2 = 0.292m" are the diameters of the master and second cylinders respectively. Expressing "F_m", obtain:


"F_m = \dfrac{F_2A_m}{A_2}\\
F_m = \dfrac{mg\cdot \pi d_m^2\cdot 4}{4\cdot \pi d_2^2} =\dfrac{mg\cdot d_m^2}{d_2^2}\\
F_m = \dfrac{2411\cdot 9.8\cdot0.021^2}{0.292^2} \approx 122N"

Answer. 122 N.